OpenStudy (anonymous):

2. Sid wants to find the height of a tree without having to climb it, but it is a cloudy day so he cannot use the shadows . he takes a mirror from his pocket and places it on thr ground 7.2m from the base of a tree. He backs up until he can see the top of the tree in the mirror, a distance of 1.2m from the mirror. If sids eyes are 1.5m above the ground what is the height of the tree?

OpenStudy (anonymous):

1.5/1.2=x/7.2 x=9